A reaction has ΔH° = 100 kJ/mol and ΔS°=250 J/mol.K. Is the reaction spontaneous at room temperature? If not, under what temperature conditions will it become spontaneous?



1
Expert's answer
2021-01-19T04:36:51-0500

"\Delta"H⁰ = 100KJ/mol and "\Delta"S⁰ = 250J/mol.K

Room temperature = 298K

"\Delta" G⁰ = "\Delta" H⁰ - T "\Delta" S⁰

= 100 KJ/mol - 298K (250J/mol.K)("\frac{1KJ}{10³J}" )

= 100 - 74.5 = 25 5 KJ

"\Delta" G⁰ = + postive , the reaction is not spontaneous under room temperature at 298 K

Now we take temperature 410K

"\Delta" G⁰ = 100KJ/mol - 410K (250J/mol.K)("\frac{1KJ}{10³J}")

= 100 - 102.5 = - 2.5 KJ

"\Delta" G⁰ = - negative , the reaction is spontaneous at 410K temperature.